# Differentially Private Relational Learning with Entity-level Privacy Guarantees

### v1

@@ 1 Introduction
Complex real-world relationships and interactions among individuals are commonly modeled as graphs, where nodes represent these individuals (or more broadly, entities) and edges depict their connections. These graph structures are invaluable for analyzing intricate systems and networks kanehisa2000kegg ; newman2003structure ; leskovec2005graphs ; leskovec2007dynamics ; kwak2010twitter ; zitnik2018modeling . Machine learning methods that leverage these structures have demonstrated significant potential in tasks reliant on such relational information, such as recommendation systems, financial network analysis, and some healthcare applications (kipf2016semi, ; hamilton2017inductive, ; bronstein2017geometric, ; stokes2020deep, ; wang2021review, ; ahmedt2021graph, ; wu2020comprehensive, ). Particularly, relational learning integrates relational information into the models, enabling them to capture and infer complex dependencies beyond isolated data points xiong2018one ; zhang2019ernie ; fang2022transferring . One promising application is fine-tuning foundation models—originally trained on text or images—using relational datasets to improve their predictive performance and adaptability to relational contexts wu2021leveraging ; zhang2022greaselm ; gao2023leveraging ; alsentzer2019publicly .
@@
Privacy concerns emerge when relational data to be incorporated into models contains sensitive information, particularly in healthcare and finance. For example, in a patient-medical record network data, patient diagnoses (entity attributes) and medical treatments (relations) represent confidential details. Machine learning models are susceptible to privacy vulnerabilities, potentially leading to unintended exposure of sensitive training data shokri2017membership ; carlini2019secret .
@@
Differential Privacy (DP) provides a principled framework to quantify and mitigate privacy risks for individual data dwork2006differential . A prominent approach to train deep learning models under the DP framework is DP-SGD song2013stochastic ; bassily2014private ; abadi2016deep ; ziller2021medical ; klause2022differentially ; fang2022differentially ; adnan2022federated . DP-SGD was initially developed for unstructured data with independent samples, where modifying a single data point impacts only one gradient term. Privacy preservation is thus achieved by controlling gradient sensitivity via per-sample gradient clipping and adding calibrated Gaussian noise to mask individual contributions and prevent privacy leakage.
@@
Moreover, DP-SGD benefits from a privacy analysis technique known as privacy amplification by subsampling abadi2016deep ; balle2018privacy ; mironov2019r ; zhu2019poission . Specifically, applying a differentially private algorithm to a randomly subsampled dataset rather than the entire dataset strengthens privacy guarantees. This mechanism aligns with mini-batch sampling, further reinforcing the privacy protection capability of DP-SGD.
@@
However, deploying DP-SGD in relational learning to protect entity-level privacy presents unique challenges due to the interconnected nature of entities and the relationship-oriented training paradigm (Figure 1). These challenges include: a) high sensitivity: entities often participate in multiple relations, thereby appearing in multiple loss terms, significantly increasing sensitivity to entity removal and the risk of sensitive information leakage; and b) coupled sampling: mini-batches in relational learning typically involve multiple interdependent sampling steps, such as sampling positive relations first and subsequently generating negative relations conditioned on these positives chen2020simple ; you2020graph . This coupled sampling approach diverges from the standard single-step independent sampling assumed by traditional DP-SGD analyses, complicating the direct application of established privacy guarantees.
@@ 2 Related Works
Differentially Private Network Analysis and Learning. Extensive research has focused on developing privacy-preserving algorithms for network analysis and learning under differential privacy (DP) guarantees (li2023private, ; mueller2022sok, )(mohamed2022differentially, ) studied differentially private community detection for stochastic block models.  (wang2013differential, ; hawkins2023node, ) considered the privatization of the graph Laplacian spectrum. (sajadmanesh2021locally, ) proposed a locally differentially private algorithm for graph neural networks (GNNs). For node-level learning tasks, (daigavane2021node, ; sajadmanesh2023gap, ; sajadmanesh2024progap, ; chien2024differentially, ; xiang2024preserving, ) developed DP methods for training GNNs. (olatunji2021releasing, ) employed a teacher-student framework to support differentially private GNN model release. Note that existing DP GNN methods are not applicable due to the distinct challenges of relational learning. These methods primarily address node label prediction tasks, with their privacy mechanisms focused on preventing leakage of node features during neighborhood aggregation, where the interaction data is used as the input features to the model. In contrast, the central privacy risk in relational learning stems from the relations themselves being part of the loss function computation, i.e., the loss format, where mini-batches are constructed through complex, coupled sampling of relations, and a single entity can participate in numerous such relations.
Privacy Amplification by Subsampling. Privacy amplification refers to an enhanced privacy guarantee that comes from subsampling the raw dataset (kasiviswanathan2011can, ; li2012sampling, ; beimel2013characterizing, ). Most existing analyses focus on neighboring datasets that differ by a single data point (balle2018privacy, ; mironov2019r, ; wang2019subsampled, ; zhu2019poission, ; steinke2022composition, ). Recently, it has been extended to group privacy, allowing neighboring datasets to differ by multiple data points (ganesh2024tight, ; schuchardt2024unified, ; jiang2024calibrating, ). These works commonly assume the dataset is an unstructured collection of independent data points, and the sampling mechanism simply outputs a subset of it. Our work instead extends the sampling mechanisms to coupled sampling, an interdependent sampling process from a composite dataset.
@@ 3 Preliminary
We use graph notations to denote the relationships between entities. Let $\mathcal{G}=(\mathcal{V},\mathcal{E},X)$ be a undirected graph defined on node (entity) set $\mathcal{V}$ and edge (relation) set $\mathcal{E}$. Node attributes $X$ assign each node $v\in\mathcal{V}$ an attribute $x_{v}$. The notion of differential privacy (DP) or Renyi DP (RDP) is defined based on neighboring unstructured dataset differed by one data point dwork2006differential ; mironov2017renyi . To extend it to graphs, neighboring graphs are defined correspondingly.
@@
DP-SGD for Private Learning. DP-SGD is a variant of the standard SGD algorithm, designed to ensure that the training process of a machine learning model adheres to differential privacy (song2013stochastic, ; abadi2016deep, ; bassily2014private, ). It assumes that the mini-batch gradient is decomposable with respect to each training sample: for a mini-batch =$\mathcal{B}=\{x_{1},\ldots,x_{b}\}$, the gradient satisfies $\mathbf{g}(\mathcal{B})=\sum_{x_{i}\in\mathcal{B}}\mathbf{g}(x_{i})$. This assumption enables bounding the sensitivity $\max_{\mathcal{B}\sim\mathcal{B}^{\prime}}\|\mathbf{g}(\mathcal{B})-\mathbf{g}(\mathcal{B}^{\prime})\|_{2}\leq C$ via per-sample gradient clipping, where each $\mathbf{g}(x_{i})$ is replaced by $\mathbf{g}(x_{i})/\max(\|\mathbf{g}(x_{i})\|,C)$. We refer to this per-sample, independently applied clipping as standard clipping throughout the paper.
Privacy Amplification by Subsampling. Let $D$ be a generic dataset, and $S$ be a sampling mechanism that produces a random mini-batch $\mathcal{B}=S(D)$. Traditionally, the dataset is a unstructured collection of samples $D=\{x_{1},...,x_{N}\}$, and $S(D)$ a random subset =$\mathcal{B}=\{x_{s_{1}},...,x_{s_{b}}\}\subseteq D$, e.g., $S$ is sampling without replacement or Poisson sampling. In relational learning, however, $D$ is a graph $\mathcal{G}$ and $S$ involves a complex, multi-stage sampling process over nodes and edges, which we will elaborate in Section 4.2. Let $M\circ S$ denote the mechanism that first applies sampling to the dataset and then runs a randomized algorithm $M$. Given a $(\varepsilon,\delta)$-DP algorithm $M$, the goal of privacy amplification is to estimate the privacy parameters $(\varepsilon^{\prime},\delta^{\prime})$ for the composed mechanism $M\circ S$ (steinke2022composition, ).
@@ 4 Problem Formulation and Main Results
Relational learning aims to develop node attribute encoders for predicting and inferring relationships between nodes (yasunaga2022linkbert, ; duan2023simteg, ; xie2023graph, ). A key advantage is that once trained, these refined encoders can be deployed for new, unseen entities to predict potential relationships in a zero-shot manner. This capability is particularly valuable in applications such as addressing zero-shot recommendation systems (leroy2010cold, ; zhuo2022tiger, ; wang2023pre, ; wu2021towards, ), and anomaly detection, where interactions involving the new entities are inherently sparse (ranshous2015anomaly, ; reiss2024zero, ). Formally, suppose a node attribute encoder $f_{\Theta}$ maps entity attributes $x_{u}$ to a feature vector $h_{u}=f_{\Theta}(x_{u})$. The goal of relational learning is to leverage the edge set $\mathcal{E}$ to learn the parameters $\Theta$. We use a scoring function $\mathrm{score}_{(u,v)}=\mathrm{score}(h_{u},h_{v})$ that reflects the likelihood of an edge existing between entities $u$ and $v$. To provide supervision signals, we use both positive (observed) edges $e^{+}\in\mathcal{E}$ and negative (missing) edges $e^{-}\notin\mathcal{E}$, typically grouped into edge tuples $T_{i}=(e^{+}_{i},e^{-}_{i,1},\ldots,e^{-}_{i,k_{\mathrm{neg}}})$ consisting of one positive edge versus $k_{\mathrm{neg}}$ negative edges (nickel2015review, ). Positive edges are usually sampled from $\mathcal{E}$, while various strategies exist for sampling negative edges (chen2020simple, ; you2020graph, ). An illustration of the mini-batch sampling process is provided in Figure 1. Once the edge tuples are constructed, training proceeds by minimizing a loss function $\mathcal{L}(\Theta,T_{i})$ that encourages higher scores for positive edges and lower scores for negative ones. Common loss functions include the InfoNCE loss, $\mathcal{L}(\Theta,T_{i})=-\log\left(\exp(\mathrm{score}_{e_{i}^{+}})/\sum_{e\in T_{i}}\exp(\mathrm{score}_{e})\right)$ (oord2018representation, ), and the Hinge loss, $\mathcal{L}(\Theta,T_{i})=\sum_{j=1}^{k}\max(0,\gamma-\mathrm{score}_{e^{+}_{i}}+\mathrm{score}_{e^{-}_{i,j}})$, with margin parameter $\gamma$ (bordes2013translating, ; wang2014knowledge, ; yang2014embedding, ; lin2015learning, ). For conceptual simplicity, our discussion focuses on binary relations where edges are either present or absent, though the approach can be extended to multi-relational settings.
Consider a mini-batch of edge tuples =$\mathcal{B}=\{T_{1},...,T_{b}\}$, whose gradient $\mathbf{g}(\mathcal{B})$ has the general form:
@@
Eq. (1) fundamentally departs from the standard DP-SGD setting, as the sensitive data unit—a node, in this case—may appear in multiple edge tuples $T_{i}$, thereby influencing several gradient terms $\mathbf{g}(T_{i})$. This setup is also technically distinct from group differential privacy (ganesh2024tight, ), since a node’s presence in positive versus negative edges can yield different contributions to the overall sensitivity. A detailed discussion on this point and sensitivity analysis, along with an adaptive gradient clipping method for sensitivity control, will be presented in Section 4.1.
@@ 4.1 Sensitivity Analysis
Let us fix a node $u^{*}$ to be removed and analyze the local sensitivity $\norm{\mathbf{g}(\mathcal{B})-\mathbf{g}(\mathcal{B})}$ where $\mathcal{B}^{\prime}$ is a neighboring mini-batch of $\mathcal{B}$ by removing a node $u^{*}$ from $\mathcal{B}$. Formally, we define neighboring mini-batches $\mathcal{B}\sim\mathcal{B}^{\prime}$ as follows: let =$\mathcal{B}=\{T_{1},\ldots,T_{b}\}$ where each edge tuple is of the form $T_{i}=(e_{i}^{+},e_{i,1}^{-},\ldots,e_{i,k_{\mathrm{neg}}}^{-})$. If node $u^{*}$ appears in the positive edge $e_{i}^{+}$, then the whole tuple $T_{i}$ is removed in $\mathcal{B}^{\prime}$; If, instead, node $u^{*}$ appears only in a negative edge $e_{i,j}^{-}$ and not in $e_{i}^{+}$, then $\mathcal{B}^{\prime}$ contains the correponding tuple $T_{i}$ but replaces all occurrences of $u^{*}$ with arbitrary nodes sampled from the graph, as shown in Fig. 1. We adopt this definition of neighboring mini-batches because it appears in our later privacy amplification analysis, corresponding to the dominating pairs that capture the statistical divergence of the subsampled mechanism when applied to neighboring graphs.
@@
For any =$\mathcal{B}=\{T_{1},\ldots,T_{b}\}$, define clipped gradient $\bar{\mathbf{g}}(T_{i})=\mathbf{g}(T_{i})/\max\{1,\norm{\mathbf{g}(T_{i})/C(T_{i},\mathcal{B})}\}$ with $C(T_{i},\mathcal{B})\coloneqq C/(\sup_{u\in T_{i}}|\mathcal{B}_{+}(u)|+|\mathcal{B}_{-}(u)|).$ Then for any pair of neighboring mini-batches $\mathcal{B}^{\prime}\sim\mathcal{B}$ that differs with respective to a node $u^{*}$, it satisfies $\norm{\bar{\mathbf{g}}(\mathcal{B})-\bar{\mathbf{g}}(\mathcal{B}^{\prime})}\leq(1+|\mathcal{B}_{-}(u^{*})|)\cdot C.$
@@
Input: Batch =$\mathcal{B}=\{T_{1},...,T_{b}\}$ where $T_{i}=(e_{i}^{+},e^{-}_{i,1},...,e^{-}_{i,k_{\mathrm{neg}}})$, gradient terms $\{\mathbf{g}(T_{1}),...,\mathbf{g}(T_{b})\}$ and gradient norm clipping threshold $C$.
@@
    $\text{freq}(v)\leftarrow\sum_{T_{i}\in\mathcal{B}}\mathbb{I}(v\in e_{i}^{+}\text{ or }v\in\bigcup_{j=1}^{k_{\mathrm{neg}}}e_{i,j}^{-})$triangleright Compute $|\mathcal{B}_{+}(v)|+|\mathcal{B}_{-}(v)|$
@@
[…]
@@
Return $\mathbf{\bar{g}}=\sum_{i}\bar{\mathbf{g}}(E_{i})$triangleright Clipped mini-batch gradient with sensitivity $C$
@@
Remark. Adaptive clipping in DP-SGD has been extensively studied for non-relational learning shulgin2024convergence , which typically adjusts the clipping threshold based on statistics of individual gradient norms, such as quantiles andrew2021differentially or distributional moments wei2025dcsgd . In contrast, our approach for relational learning employs a normalization strategy based on node occurrences within a batch to address the unique challenges of this setting.

### v3

@@ 1 Introduction
Complex real-world relationships and interactions among individuals are commonly modeled as graphs, where nodes represent these individuals (or more broadly, entities) and edges depict their connections. These graph structures are invaluable for analyzing intricate systems and networks Kanehisa and Goto (2000); Newman (2003); Leskovec et al. (2005, 2007); Kwak et al. (2010); Zitnik et al. (2018). Machine learning methods that leverage these structures have demonstrated significant potential in tasks reliant on such relational information, such as recommendation systems, financial network analysis, and some healthcare applications (Kipf and Welling, 2017; Hamilton et al., 2017; Bronstein et al., 2017; Stokes et al., 2020; Wang et al., 2021; Ahmedt-Aristizabal et al., 2021; Wu et al., 2020). Particularly, relational learning integrates relational information into the models, enabling them to capture and infer complex dependencies beyond isolated data points Xiong et al. (2018); Zhang et al. (2019); Fang et al. (2022a). One promising application is fine-tuning foundation models—originally trained on text or images—using relational datasets to improve their predictive performance and adaptability to relational contexts Wu et al. (2021b); Zhang et al. (2022); Gao et al. (2023); Alsentzer et al. (2019).
@@
Privacy concerns emerge when relational data to be incorporated into models contains sensitive information, particularly in healthcare and finance. For example, in a patient-medical record network data, patient diagnoses (entity attributes) and medical treatments (relations) represent confidential details. Machine learning models are susceptible to privacy vulnerabilities, potentially leading to unintended exposure of sensitive training data Shokri et al. (2017); Carlini et al. (2019).
@@
Differential Privacy (DP) provides a principled framework to quantify and mitigate privacy risks for individual data Dwork (2006). A prominent approach to train deep learning models under the DP framework is DP-SGD Song et al. (2013); Bassily et al. (2014); Abadi et al. (2016); Ziller et al. (2021); Klause et al. (2022); Fang et al. (2022b); Adnan et al. (2022). DP-SGD was initially developed for unstructured data with independent samples, where modifying a single data point impacts only one gradient term. Privacy preservation is thus achieved by controlling gradient sensitivity via per-sample gradient clipping and adding calibrated Gaussian noise to mask individual contributions and prevent privacy leakage.
@@
Moreover, DP-SGD benefits from a privacy analysis technique known as privacy amplification by subsampling Abadi et al. (2016); Balle et al. (2018); Mironov et al. (2019); Zhu and Wang (2019). Specifically, applying a differentially private algorithm to a randomly subsampled dataset rather than the entire dataset strengthens privacy guarantees. This mechanism aligns with mini-batch sampling, further reinforcing the privacy protection capability of DP-SGD.
@@
However, deploying DP-SGD in relational learning to protect entity-level privacy presents unique challenges due to the interconnected nature of entities and the relationship-oriented training paradigm (Figure 1). These challenges include: a) high sensitivity: entities often participate in multiple relations, thereby appearing in multiple loss terms, significantly increasing sensitivity to entity removal and the risk of sensitive information leakage; and b) coupled sampling: mini-batches in relational learning typically involve multiple interdependent sampling steps, such as sampling positive relations first and subsequently generating negative relations conditioned on these positives Chen et al. (2020); You et al. (2020). This coupled sampling approach diverges from the standard single-step independent sampling assumed by traditional DP-SGD analyses, complicating the direct application of established privacy guarantees.
@@ 2 Related Works
Differentially Private Network Analysis and Learning. Extensive research has focused on developing privacy-preserving algorithms for network analysis and learning under differential privacy (DP) guarantees (Li et al., 2023; Mueller et al., 2022)(Mohamed et al., 2022) studied differentially private community detection for stochastic block models.  (Wang et al., 2013; Hawkins et al., 2023) considered the privatization of the graph Laplacian spectrum. (Sajadmanesh and Gatica-Perez, 2021) proposed a locally differentially private algorithm for graph neural networks (GNNs). For node-level learning tasks, (Daigavane et al., 2021; Sajadmanesh et al., 2023; Sajadmanesh and Gatica-Perez, 2024; Chien et al., 2024; Xiang et al., 2024) developed DP methods for training GNNs. (Olatunji et al., 2023) employed a teacher-student framework to support differentially private GNN model release. Note that existing DP GNN methods are not applicable due to the distinct challenges of relational learning. These methods primarily address node label prediction tasks, with their privacy mechanisms focused on preventing leakage of node features during neighborhood aggregation, where the interaction data is used as the input features to the model. In contrast, the central privacy risk in relational learning stems from the relations themselves being part of the loss function computation, i.e., the loss format, where mini-batches are constructed through complex, coupled sampling of relations, and a single entity can participate in numerous such relations.
Privacy Amplification by Subsampling. Privacy amplification refers to an enhanced privacy guarantee that comes from subsampling the raw dataset (Kasiviswanathan et al., 2011; Li et al., 2012; Beimel et al., 2013). Most existing analyses focus on neighboring datasets that differ by a single data point (Balle et al., 2018; Mironov et al., 2019; Wang et al., 2019; Zhu and Wang, 2019; Steinke, 2022). Recently, it has been extended to group privacy, allowing neighboring datasets to differ by multiple data points (Ganesh, 2024; Schuchardt et al., 2024; Jiang et al., 2024). These works commonly assume the dataset is an unstructured collection of independent data points, and the sampling mechanism simply outputs a subset of it. Our work instead extends the sampling mechanisms to coupled sampling, an interdependent sampling process from a composite dataset.
@@ 3 Preliminary
We use graph notations to denote the relationships between entities. Let $\mathcal{G}=(\mathcal{V},\mathcal{E},X)$ be a undirected graph defined on node (entity) set $\mathcal{V}$ and edge (relation) set $\mathcal{E}$. Node attributes $X$ assign each node $v\in\mathcal{V}$ an attribute $x_{v}$. The notion of differential privacy (DP) or Renyi DP (RDP) is defined based on neighboring unstructured dataset differed by one data point Dwork (2006); Mironov (2017). To extend it to graphs, neighboring graphs are defined correspondingly.
@@
DP-SGD for Private Learning. DP-SGD is a variant of the standard SGD algorithm, designed to ensure that the training process of a machine learning model adheres to differential privacy (Song et al., 2013; Abadi et al., 2016; Bassily et al., 2014). It assumes that the mini-batch gradient is decomposable with respect to each training sample: for a mini-batch $\mathcal{B}=\{x_{1},\ldots,x_{b}\}$, the gradient satisfies $\mathbf{g}(\mathcal{B})=\sum_{x_{i}\in\mathcal{B}}\mathbf{g}(x_{i})$. This assumption enables bounding the sensitivity $\max_{\mathcal{B}\sim\mathcal{B}^{\prime}}\|\mathbf{g}(\mathcal{B})-\mathbf{g}(\mathcal{B}^{\prime})\|_{2}\leq C$ via per-sample gradient clipping, where each $\mathbf{g}(x_{i})$ is replaced by $\mathbf{g}(x_{i})/\max(\|\mathbf{g}(x_{i})\|,C)$. We refer to this per-sample, independently applied clipping as standard clipping throughout the paper.
Privacy Amplification by Subsampling. Let $D$ be a generic dataset, and $S$ be a sampling mechanism that produces a random mini-batch $\mathcal{B}=S(D)$. Traditionally, the dataset is a unstructured collection of samples $D=\{x_{1},...,x_{N}\}$, and $S(D)$ a random subset $\mathcal{B}=\{x_{s_{1}},...,x_{s_{b}}\}\subseteq D$, e.g., $S$ is sampling without replacement or Poisson sampling. In relational learning, however, $D$ is a graph $\mathcal{G}$ and $S$ involves a complex, multi-stage sampling process over nodes and edges, which we will elaborate in Section 4.2. Let $M\circ S$ denote the mechanism that first applies sampling to the dataset and then runs a randomized algorithm $M$. Given a $(\varepsilon,\delta)$-DP algorithm $M$, the goal of privacy amplification is to estimate the privacy parameters $(\varepsilon^{\prime},\delta^{\prime})$ for the composed mechanism $M\circ S$ (Steinke, 2022).
@@ 4 Problem Formulation and Main Results
Relational learning aims to develop node attribute encoders for predicting and inferring relationships between nodes (Yasunaga et al., 2022; Duan et al., 2023; Xie et al., 2023). A key advantage is that once trained, these refined encoders can be deployed for new, unseen entities to predict potential relationships in a zero-shot manner. This capability is particularly valuable in applications such as addressing zero-shot recommendation systems (Leroy et al., 2010; Zhuo et al., 2022; Wang et al., 2023; Wu et al., 2021a), and anomaly detection, where interactions involving the new entities are inherently sparse (Ranshous et al., 2015; Reiss et al., 2024). Formally, suppose a node attribute encoder $f_{\Theta}$ maps entity attributes $x_{u}$ to a feature vector $h_{u}=f_{\Theta}(x_{u})$. The goal of relational learning is to leverage the edge set $\mathcal{E}$ to learn the parameters $\Theta$. We use a scoring function $\mathrm{score}_{(u,v)}=\mathrm{score}(h_{u},h_{v})$ that reflects the likelihood of an edge existing between entities $u$ and $v$. To provide supervision signals, we use both positive (observed) edges $e^{+}\in\mathcal{E}$ and negative (missing) edges $e^{-}\notin\mathcal{E}$, typically grouped into edge tuples $T_{i}=(e^{+}_{i},e^{-}_{i,1},\ldots,e^{-}_{i,k_{\mathrm{neg}}})$ consisting of one positive edge versus $k_{\mathrm{neg}}$ negative edges (Nickel et al., 2015). Positive edges are usually sampled from $\mathcal{E}$, while various strategies exist for sampling negative edges (Chen et al., 2020; You et al., 2020). An illustration of the mini-batch sampling process is provided in Figure 1. Once the edge tuples are constructed, training proceeds by minimizing a loss function $\mathcal{L}(\Theta,T_{i})$ that encourages higher scores for positive edges and lower scores for negative ones. Common loss functions include the InfoNCE loss, $\mathcal{L}(\Theta,T_{i})=-\log\left(\exp(\mathrm{score}_{e_{i}^{+}})/\sum_{e\in T_{i}}\exp(\mathrm{score}_{e})\right)$ (Oord et al., 2018), and the Hinge loss, $\mathcal{L}(\Theta,T_{i})=\sum_{j=1}^{k}\max(0,\gamma-\mathrm{score}_{e^{+}_{i}}+\mathrm{score}_{e^{-}_{i,j}})$, with margin parameter $\gamma$ (Bordes et al., 2013; Wang et al., 2014; Yang et al., 2015; Lin et al., 2015). For conceptual simplicity, our discussion focuses on binary relations where edges are either present or absent, though the approach can be extended to multi-relational settings.
Consider a mini-batch of edge tuples $\mathcal{B}=\{T_{1},...,T_{b}\}$, whose gradient $\mathbf{g}(\mathcal{B})$ has the general form:
@@
Eq. (1) fundamentally departs from the standard DP-SGD setting, as the sensitive data unit—a node, in this case—may appear in multiple edge tuples $T_{i}$, thereby influencing several gradient terms $\mathbf{g}(T_{i})$. This setup is also technically distinct from group differential privacy (Ganesh, 2024), since a node’s presence in positive versus negative edges can yield different contributions to the overall sensitivity. A detailed discussion on this point and sensitivity analysis, along with an adaptive gradient clipping method for sensitivity control, will be presented in Section 4.1.
@@ 4.1 Sensitivity Analysis
Let us fix a node $u^{*}$ to be removed and analyze the local sensitivity $\norm{\mathbf{g}(\mathcal{B})-\mathbf{g}(\mathcal{B})}$ where $\mathcal{B}^{\prime}$ is a neighboring mini-batch of $\mathcal{B}$ by removing a node $u^{*}$ from $\mathcal{B}$. Formally, we define neighboring mini-batches $\mathcal{B}\sim\mathcal{B}^{\prime}$ as follows: let $\mathcal{B}=\{T_{1},\ldots,T_{b}\}$ where each edge tuple is of the form $T_{i}=(e_{i}^{+},e_{i,1}^{-},\ldots,e_{i,k_{\mathrm{neg}}}^{-})$. If node $u^{*}$ appears in the positive edge $e_{i}^{+}$, then the whole tuple $T_{i}$ is removed in $\mathcal{B}^{\prime}$; If, instead, node $u^{*}$ appears only in a negative edge $e_{i,j}^{-}$ and not in $e_{i}^{+}$, then $\mathcal{B}^{\prime}$ contains the correponding tuple $T_{i}$ but replaces all occurrences of $u^{*}$ with arbitrary nodes sampled from the graph, as shown in Fig. 1. We adopt this definition of neighboring mini-batches because it appears in our later privacy amplification analysis, corresponding to the dominating pairs that capture the statistical divergence of the subsampled mechanism when applied to neighboring graphs.
@@
For any $\mathcal{B}=\{T_{1},\ldots,T_{b}\}$, define clipped gradient $\bar{\mathbf{g}}(T_{i})=\mathbf{g}(T_{i})/\max\{1,\norm{\mathbf{g}(T_{i})/C(T_{i},\mathcal{B})}\}$ with $C(T_{i},\mathcal{B})\coloneqq C/(\sup_{u\in T_{i}}|\mathcal{B}_{+}(u)|+|\mathcal{B}_{-}(u)|).$ Then for any pair of neighboring mini-batches $\mathcal{B}^{\prime}\sim\mathcal{B}$ that differs with respective to a node $u^{*}$, it satisfies $\norm{\bar{\mathbf{g}}(\mathcal{B})-\bar{\mathbf{g}}(\mathcal{B}^{\prime})}\leq(1+|\mathcal{B}_{-}(u^{*})|)\cdot C.$
@@
Input: Batch $\mathcal{B}=\{T_{1},...,T_{b}\}$ where $T_{i}=(e_{i}^{+},e^{-}_{i,1},...,e^{-}_{i,k_{\mathrm{neg}}})$, gradient terms $\{\mathbf{g}(T_{1}),...,\mathbf{g}(T_{b})\}$ and gradient norm clipping threshold $C$.
@@
  $\text{freq}(v)\leftarrow\sum_{T_{i}\in\mathcal{B}}\mathbb{I}(v\in e_{i}^{+}\text{ or }v\in\bigcup_{j=1}^{k_{\mathrm{neg}}}e_{i,j}^{-})$ $\triangleright$ Compute $|\mathcal{B}_{+}(v)|+|\mathcal{B}_{-}(v)|$
@@
  $\text{max-freq}(T_{i})\leftarrow\max_{v\in\mathcal{V}_{T_{i}}}\text{freq}(v)$
@@
Return $\mathbf{\bar{g}}=\sum_{i}\bar{\mathbf{g}}(E_{i})$ $\triangleright$ Clipped mini-batch gradient with sensitivity $C$
@@
Remark. Adaptive clipping in DP-SGD has been extensively studied for non-relational learning Shulgin and Richtárik (2024), which typically adjusts the clipping threshold based on statistics of individual gradient norms, such as quantiles Andrew et al. (2021) or distributional moments Wei et al. (2025). In contrast, our approach for relational learning employs a normalization strategy based on node occurrences within a batch to address the unique challenges of this setting.